Window Help Question 13 474 pts The ionization constant of water at a temperature above 25°C is 3.5 10-14. What is the pH of show work
Answer #1

Kwa [Hzot] [ot] [H3ot) = [04] = x x² = 3.5x10-14 x = 1.870 8 2 864 x 10 M = (Hgot] pa-log (1304) = -lag (1-87682869 x 10) =